### v1

@@ 5 Object Detection in Autonomous Vehicles
The MonoTAKD [138] framework introduces a three-part distillation strategy to enhance monocular 3D detection, comprising a LiDAR-based teacher, a camera-based teaching assistant (TA), and a camera-based student. Using a ResNet-50 backbone, the TA network integrates accurate ground-truth depth maps with visual features to produce high-fidelity BEV representations, referred to as 3D visual knowledge. These representations guide the student via intra-modal distillation, mitigating errors caused by monocular depth estimation. The student, architecturally aligned with the TA but relying on predicted depth, is structured with two feature branches: one dedicated to learning visual cues from the TA, and another designed to absorb spatial information from residual BEV features obtained by subtracting the TA’s output from that of the LiDAR-based teacher. Cross-modal residual distillation transfers only these spatial cues, narrowing the modality gap without requiring the student to replicate the entire LiDAR feature space. A Spatial Alignment Module, combining dilated and deformable convolutions with channel attention, further refines spatial branch outputs. At the same time, a Feature Fusion Module merges the visual and spatial streams for final detection. This targeted multi-stage guidance enables the student to capture a richer geometric context than conventional monocular models.
@@
[137] MonoDGP 2025 CVPR Res50 Mono 35.24 25.23 22.02 26.35 18.72 15.97 - - - - - -
@@
[138] MonoTAKD 2025 CVPR Res50 Mono 38.75 27.76 24.14 27.91 19.43 16.51 16.15 10.41 9.68 13.54 7.23 6.86
@@
[140] FD3D 2024 AAAI MonoDLE Mono 34.20 23.72 20.76 25.38 17.12 14.50 - - - - - -
@@
[141] MonoCD 2024 CVPR DLA-34 Mono 33.41 22.81 19.57 25.53 16.59 14.53 - - - - - -
@@
[139] MonoDiff 2024 CVPR Res50-FPN Mono - - - 30.18 21.02 18.16 13.51 8.94 7.28 8.52 4.35 3.78
@@
[142] GUPNet++ 2024 TPAMI DLA-34 Depth - - - 24.99 16.48 14.58 12.45 8.13 6.91 6.71 3.91 3.80
@@
[143] MonoDETR 2023 ICCV Res50 Mono 33.60 22.11 18.60 25.00 16.47 13.58 - - - - - -
@@
[144] MonoUNI 2023 NeurIPS Res50 Mono - - - 24.75 16.73 13.49 15.78 1.34 8.74 7.34 4.28 3.78
@@
[145] DEVIANT 2023 CVPR Res50 Mono 29.65 20.44 17.43 21.88 14.46 11.89 - - - - - -
@@
[146] MonoPGC 2023 CVPR DLA34 Mono 32.50 23.14 20.30 24.68 17.17 14.14 - - - - - -
@@
[147] MonoATT 2023 CVPR DLA-34 Mono 36.87 24.42 21.88 24.72 17.37 15.00 10.55 6.66 5.43 5.74 3.68 2.94
@@
[149] MonoDDE 2022 CVPR DLA34 Mono 33.58 23.46 20.37 24.93 17.14 15.10 11.13 7.32 6.67 5.94 3.78 3.33
@@
[150] MonoGround 2022 CVPR DLA-34 Mono 30.07 20.47 17.74 21.37 14.36 12.62 12.37 7.89 7.13 4.62 2.68 2.53
@@
[151] MonoJSG 2022 CVPR DLA-34 Mono 32.59 21.26 18.18 24.69 16.14 13.64 11.94 7.36 6.03 8.03 3.87 3.33
@@
[152] Kinematic3D 2022 ECCV Dense121 Mono 26.69 17.52 13.10 19.07 12.72 9.17 - - - - - -
@@
[153] ImVoxelNet 2022 WACV Res50 Mono - - - 17.15 10.97 9.15 - - - - - -
@@
[154] ImVoxelNet 2022 CVPR DLA34 Mono - - - 20.10 12.99 10.50 12.47 7.62 6.72 1.52 0.85 0.94
@@
[154] MonoFlex 2022 CVPR DLA34 Mono - - - 21.75 14.94 13.07 11.87 7.66 6.82 5.48 3.50 2.99
@@
[155] CMKD 2022 ECCV Res50 Depth - - - 25.09 16.99 15.30 17.79 11.69 10.09 9.60 5.24 4.50
@@
[157] GUPNet 2021 ICCV DLA34 Mono - - - 20.11 14.20 11.77 14.72 9.53 7.87 4.18 2.65 2.09
@@
[158] MonoDLE 2021 CVPR DLA34 Mono 24.79 18.89 16.00 17.23 12.26 10.29 5.34 3.28 2.83 4.59 2.66 2.45
@@
[159] MonoRCNN 2021 ICCV Res50 Mono 25.48 18.11 14.10 18.36 12.65 10.03 11.21 7.28 5.85 2.89 1.67 1.54
@@
[160] MonoFlex 2021 CVPR DLA34 Mono 28.23 19.75 16.89 19.94 13.89 12.07 9.43 6.31 5.26 4.17 2.35 2.04
@@
[161] MonoRUn 2021 CVPR Res101 Mono 27.94 17.34 15.24 19.65 12.30 10.58 10.88 6.78 5.83 1.01 0.61 0.48
@@
[162] DDMP-3D 2021 CVPR Res101 Depth 28.08 17.89 13.44 19.71 12.78 9.80 4.93 3.55 3.01 4.18 2.35 2.04
@@
[163] MonoPair 2020 CVPR DLA34 Mono - - - 13.04 9.99 8.65 10.02 6.68 5.53 3.79 2.12 1.83
@@
[166] MoVi-3D 2020 ECCV Res34 Mono - - - 15.19 10.90 9.26 8.99 5.44 4.57 1.08 0.63 0.70
@@
[167] DA-3Ddet 2020 ECCV Res101 Mono - - - 16.77 11.50 8.93 8.7 7.1 6.7 14.5 11.5 11.5
@@
[168] MonoDIS 2019 ICCV Res34 Mono 17.23 13.19 11.12 10.37 7.94 6.40 - - - - - -
@@
[136] M3D-RPN 2019 ICCV Dense121 Mono - - - 14.76 9.71 7.42 4.92 3.48 2.94 0.94 0.65 0.47
@@
Table 16 presents a comparative analysis of 3D object detection methods using camera-based data from the NuScenes dataset, where the input is 2D images and the output is 3D bounding boxes. From Table 16, we provide a detailed discussion of two selected models: DETR3D [169], recognized as a well-known and widely adopted model in this field; and CorrBEV [170], which achieved the highest results on the NuScenes dataset. It should be noted that the values highlighted in red represent the best performance achieved for each evaluation metric across all listed methods.
DETR3D [169] addresses multi-view 3D object detection by establishing a direct mapping between a sparse set of learned 3D object queries and multi-camera image features, bypassing any intermediate dense geometric reconstruction. Feature extraction is performed by a shared ResNet–FPN backbone, producing multi-scale image representations for each view. Each object query contains a latent 3D hypothesis, decoded into a reference point in the scene. Through calibrated camera intrinsics and extrinsics, these points are back-projected into each image, where bilinear interpolation retrieves the corresponding feature vectors from all pyramid levels. Aggregated features are merged into the queries and fine-tuned using multi-head self-attention. This projection–refinement cycle repeats across multiple transformer decoder layers, with each stage producing updated classifications and 3D box regressions. DETR3D maintains an end-to-end pipeline that improves efficiency without sacrificing competitive detection accuracy by avoiding dense depth estimation, voxelization, and non-maximum suppression.
@@
[170] CorrBEV-fm 2025 CVPR Res101 Muiti-View 41.30 50.70 0.646 0.265 0.482 0.473 0.134
@@
[170] CorrBEV-sp 2025 CVPR Res101 Muiti-View 56.20 64.00 0.488 0.241 0.303 0.243 0.123
@@
[171] RoPETR 2025 ArXiv V2-99 Muiti-View 52.90 61.40 0.537 0.255 0.289 0.229 0.195
@@
[172] BEVFormer 2024 IEEE Res101 Muiti-View 44.50 53.50 0.631 0.257 0.405 0.435 0.143
@@
[173] RayDN 2024 ECCV Res50 Muiti-View 46.90 56.30 0.579 0.264 0.433 0.256 0.187
@@
[174] StreamPETR 2023 ICCV Res50 Muiti-View 45.00 55.00 0.569 0.267 0.413 0.265 0.196
@@
[175] PETRv2 2023 ICCV Res50 Muiti-View 34.90 45.60 0.700 0.275 0.580 0.437 0.187
@@
[176] BEVDepth 2023 AAAI Res50 Muiti-View 35.10 47.50 0.639 0.267 0.479 0.428 0.198
@@
[177] SparseBEV 2023 ICCV Res50 Muiti-View 44.80 55.80 0.581 0.271 0.373 0.247 0.190
@@
[178] HoP 2023 ICCV Res50 Muiti-View 40.00 51.00 0.607 0.275 0.515 0.286 0.216
@@
[179] PolarFormer 2023 AAAI Res101 Muiti-View 41.50 47.00 0.657 0.263 0.405 0.911 0.139
@@
[180] FrustumFormer 2023 CVPR Res101 Muiti-View 47.80 56.10 0.575 0.257 0.402 0.411 0.132
@@
[181] FB-BEV 2023 ICCV V2-99 Muiti-View 53.70 62.40 0.439 0.250 0.358 0.270 0.128
@@
[182] MV2D 2023 ICCV V2-99 Muiti-View 46.30 51.40 0.542 0.247 0.403 0.854 0.127
@@
[183] CAPE 2023 CVPR V2-99 Muiti-View 52.50 61.00 0.503 0.242 0.361 0.306 0.114
@@
[184] 3DPPE 2023 ICCV VoV-99 Muiti-View 51.40 56.60 0.569 0.255 0.394 0.796 0.138
@@
[185] PETR 2022 ECCV Res101 Muiti-View 37.00 45.50 0.647 0.251 0.433 0.933 0.143
@@
[186] UVTR 2022 NeurIPS Res101 Muiti-View 37.90 48.30 0.731 0.267 0.350 0.510 0.200
@@
[187] BEVStereo 2022 AAAI Res50 Muiti-View 37.20 50.00 0.597 0.270 0.438 0.367 0.190
@@
[188] PGD 2022 CoRL Res101 Muiti-View 38.60 44.80 0.626 0.245 0.451 1.509 0.127
@@
[169] DETR3D 2022 CoRL Res101+FPN Muiti-View 41.20 47.90 0.641 0.255 0.394 0.845 0.133
@@
[189] Graph-DETR3D 2022 ACM Res101 Muiti-View 41.80 47.20 0.668 0.250 0.440 0.876 0.139
@@
[190] FCOS3D 2021 ICCV Res101 Muiti-View 35.80 42.80 0.690 0.249 0.452 1.434 0.124
@@
CorrBEV [170] improves upon transformer-based multi-view detection frameworks like BEVFormer. The task is completed by integrating a correlation-driven multi-modal approach to counteract the loss of discriminative information under occlusion. Its ”Multi-modal Prototype Generator” forms category-specific visual prototypes from cropped ground-truth object patches and derives complementary language prototypes from class names processed through a lightweight BERT encoder. These visual–linguistic embeddings are combined and introduced into the backbone features using a depth-wise correlation mechanism. The Correlation-guided Query Learner leverages these correlation features in two complementary ways: first, to initialize object queries using high-confidence spatial locations, improving recall for partially occluded objects; and second, to perform dual-path mixed sampling in which 3D queries aggregate both backbone and correlation-derived features for richer semantic–geometric representation. An occlusion-aware trainer further augments the system via pseudo-occlusion, randomly masking pixels in high-visibility samples, then a contrastive alignment loss that brings occluded and obvious instances of the same category closer in the embedding space. This modular design yields consistent gains across all occlusion levels and integrates seamlessly into diverse baseline architectures with minimal computational cost.
@@
Table 17 summarizes YOLO-based methods evaluated on the camera-based KITTI dataset for 2D object detection. Monocular 2D object detection methods have evolved to advance efficiency, representational depth, and domain-specific tuning. Early designs, such as YOLOV3-tiny [191], prioritized minimizing computation cost to improve efficiency with a precision trade-off. Later variants, including YOLOv5n [192] with its scaled CSP-based backbone (C2f) derived from CSPDarknet53 and YOLOv6n [193] employing the EfficientRep backbone, coupled backbone revisions with refined neck structures and decoupled heads to raise accuracy without undermining real-time operation. YOLOv7 [194] extended these gains by introducing re-parameterized convolution, a broadened trainable bag-of-freebies, and coarse-to-fine hierarchical label assignment. In turn, YOLOv8n [192] replaced C3 modules with C2f blocks. It adopted an anchor-free head, simplifying the pipeline while preserving detection strength, whereas YOLOv8-Lite [195] targeted resource-limited environments through compact backbone choices, lightweight feature aggregation, and attention mechanisms. ShuffYOLOX [196] modified the YOLOX framework by utilizing the lightweight ShuffDet backbone and integrating ECA attention within the PAFPN, which reduces architectural complexity without negatively affecting detection capabilities.
@@
Domain-specific optimization has emerged as a critical trend within recent iterations. YOLO-Vehicle-v1s [197] was designed for vehicle detection in adverse visual conditions, incorporating image dehazing and multimodal image-text fusion to enhance robustness. YOLOv8-RTDAV [192] targeted road-target detection by embedding ECA-based C2f modules, DySample-based upsampling, and EIoU loss, yielding notable improvements in small-object accuracy. Transformer-CNN hybridization appeared in SwinYOLOv5s [198], which leveraged swin transformer attention with adaptive self-concat fusion to capture global context while mitigating false detections. Similarly, improved YOLOv5 [199] employed NAS-pruned branches, coordinate attention, and structural re-parameterization to achieve both speed and high precision for small-object detection. The latest designs, YOLOv9t [200] and YOLOv10n [201], represent a shift toward more fundamental architectural changes: the introduction of the GELAN backbone and programmable gradient optimization in YOLOv9t improves training stability, while YOLOv10n adopts an NMS-free consistent dual-assignment strategy to harmonize efficiency with accuracy. Collectively, these methods illustrate a clear trajectory from minimalist, speed-oriented architectures to hybrid designs that balance inference efficiency with enhanced feature representation, addressing the increasingly diverse deployment requirements of modern detection systems.
@@
[198] SwinYOLOv5s 2025 MDPI MobileNetV3 Mono 96.00% 90.20% 95.70% 50.30M - -
@@
[192] YOLOv8-RTDAV 2025 Scientific Reports C2f Mono 86.90% 81.10% 88.00% 2.81M 5.78MB 102.00

### v2

@@ 5 Object Detection in Autonomous Vehicles
The MonoTAKD [Liu_2025_CVPR] framework introduces a three-part distillation strategy to enhance monocular 3D detection, comprising a LiDAR-based teacher, a camera-based teaching assistant (TA), and a camera-based student. Using a ResNet-50 backbone, the TA network integrates accurate ground-truth depth maps with visual features to produce high-fidelity BEV representations, referred to as 3D visual knowledge. These representations guide the student via intra-modal distillation, mitigating errors caused by monocular depth estimation. The student, architecturally aligned with the TA but relying on predicted depth, is structured with two feature branches: one dedicated to learning visual cues from the TA, and another designed to absorb spatial information from residual BEV features obtained by subtracting the TA’s output from that of the LiDAR-based teacher. Cross-modal residual distillation transfers only these spatial cues, narrowing the modality gap without requiring the student to replicate the entire LiDAR feature space. A Spatial Alignment Module, combining dilated and deformable convolutions with channel attention, further refines spatial branch outputs. At the same time, a Feature Fusion Module merges the visual and spatial streams for final detection. This targeted multi-stage guidance enables the student to capture a richer geometric context than conventional monocular models.
@@
[Pu_2025_CVPR] MonoDGP 2025 CVPR Res50 Mono 35.24 25.23 22.02 26.35 18.72 15.97 - - - - - -
@@
[Liu_2025_CVPR] MonoTAKD 2025 CVPR Res50 Mono 38.75 27.76 24.14 27.91 19.43 16.51 16.15 10.41 9.68 13.54 7.23 6.86
@@
[wu2024fd3d] FD3D 2024 AAAI MonoDLE Mono 34.20 23.72 20.76 25.38 17.12 14.50 - - - - - -
@@
[Yan_2024_CVPR] MonoCD 2024 CVPR DLA-34 Mono 33.41 22.81 19.57 25.53 16.59 14.53 - - - - - -
@@
[Ranasinghe_2024_CVPR] MonoDiff 2024 CVPR Res50-FPN Mono - - - 30.18 21.02 18.16 13.51 8.94 7.28 8.52 4.35 3.78
@@
[lu2024gupnet++] GUPNet++ 2024 TPAMI DLA-34 Depth - - - 24.99 16.48 14.58 12.45 8.13 6.91 6.71 3.91 3.80
@@
[Zhang_2023_ICCV] MonoDETR 2023 ICCV Res50 Mono 33.60 22.11 18.60 25.00 16.47 13.58 - - - - - -
@@
[jinrang2023monouni] MonoUNI 2023 NeurIPS Res50 Mono - - - 24.75 16.73 13.49 15.78 1.34 8.74 7.34 4.28 3.78
@@
[kumar2022deviant] DEVIANT 2023 CVPR Res50 Mono 29.65 20.44 17.43 21.88 14.46 11.89 - - - - - -
@@
[wu2023monopgc] MonoPGC 2023 CVPR DLA34 Mono 32.50 23.14 20.30 24.68 17.17 14.14 - - - - - -
@@
[Zhou_2023_CVPR] MonoATT 2023 CVPR DLA-34 Mono 36.87 24.42 21.88 24.72 17.37 15.00 10.55 6.66 5.43 5.74 3.68 2.94
@@
[Li_2022_CVPR] MonoDDE 2022 CVPR DLA34 Mono 33.58 23.46 20.37 24.93 17.14 15.10 11.13 7.32 6.67 5.94 3.78 3.33
@@
[Qin_2022_CVPR] MonoGround 2022 CVPR DLA-34 Mono 30.07 20.47 17.74 21.37 14.36 12.62 12.37 7.89 7.13 4.62 2.68 2.53
@@
[Lian_2022_CVPR] MonoJSG 2022 CVPR DLA-34 Mono 32.59 21.26 18.18 24.69 16.14 13.64 11.94 7.36 6.03 8.03 3.87 3.33
@@
[brazil2020kinematic] Kinematic3D 2022 ECCV Dense121 Mono 26.69 17.52 13.10 19.07 12.72 9.17 - - - - - -
@@
[Rukhovich_2022_WACV] ImVoxelNet 2022 WACV Res50 Mono - - - 17.15 10.97 9.15 - - - - - -
@@
[Gu_2022_CVPR] ImVoxelNet 2022 CVPR DLA34 Mono - - - 20.10 12.99 10.50 12.47 7.62 6.72 1.52 0.85 0.94
@@
[Gu_2022_CVPR] MonoFlex 2022 CVPR DLA34 Mono - - - 21.75 14.94 13.07 11.87 7.66 6.82 5.48 3.50 2.99
@@
[hong2022cross] CMKD 2022 ECCV Res50 Depth - - - 25.09 16.99 15.30 17.79 11.69 10.09 9.60 5.24 4.50
@@
[Lu_2021_ICCV] GUPNet 2021 ICCV DLA34 Mono - - - 20.11 14.20 11.77 14.72 9.53 7.87 4.18 2.65 2.09
@@
[Ma_2021_CVPR] MonoDLE 2021 CVPR DLA34 Mono 24.79 18.89 16.00 17.23 12.26 10.29 5.34 3.28 2.83 4.59 2.66 2.45
@@
[Shi_2021_ICCV] MonoRCNN 2021 ICCV Res50 Mono 25.48 18.11 14.10 18.36 12.65 10.03 11.21 7.28 5.85 2.89 1.67 1.54
@@
[Zhang_2021_CVPR] MonoFlex 2021 CVPR DLA34 Mono 28.23 19.75 16.89 19.94 13.89 12.07 9.43 6.31 5.26 4.17 2.35 2.04
@@
[Chen_2021_CVPR] MonoRUn 2021 CVPR Res101 Mono 27.94 17.34 15.24 19.65 12.30 10.58 10.88 6.78 5.83 1.01 0.61 0.48
@@
[Wang_2021_CVPR] DDMP-3D 2021 CVPR Res101 Depth 28.08 17.89 13.44 19.71 12.78 9.80 4.93 3.55 3.01 4.18 2.35 2.04
@@
[chen2020monopair] MonoPair 2020 CVPR DLA34 Mono - - - 13.04 9.99 8.65 10.02 6.68 5.53 3.79 2.12 1.83
@@
[simonelli2020towards] MoVi-3D 2020 ECCV Res34 Mono - - - 15.19 10.90 9.26 8.99 5.44 4.57 1.08 0.63 0.70
@@
[ye2020monocular] DA-3Ddet 2020 ECCV Res101 Mono - - - 16.77 11.50 8.93 8.7 7.1 6.7 14.5 11.5 11.5
@@
[simonelli2019disentangling] MonoDIS 2019 ICCV Res34 Mono 17.23 13.19 11.12 10.37 7.94 6.40 - - - - - -
@@
[brazil2019m3d] M3D-RPN 2019 ICCV Dense121 Mono - - - 14.76 9.71 7.42 4.92 3.48 2.94 0.94 0.65 0.47
@@
Table 16 presents a comparative analysis of 3D object detection methods using camera-based data from the NuScenes dataset, where the input is 2D images and the output is 3D bounding boxes. From Table 16, we provide a detailed discussion of two selected models: DETR3D [wang2022detr3d], recognized as a well-known and widely adopted model in this field; and CorrBEV [Xue_2025_CVPR], which achieved the highest results on the NuScenes dataset. It should be noted that the values highlighted in red represent the best performance achieved for each evaluation metric across all listed methods.
DETR3D [wang2022detr3d] addresses multi-view 3D object detection by establishing a direct mapping between a sparse set of learned 3D object queries and multi-camera image features, bypassing any intermediate dense geometric reconstruction. Feature extraction is performed by a shared ResNet–FPN backbone, producing multi-scale image representations for each view. Each object query contains a latent 3D hypothesis, decoded into a reference point in the scene. Through calibrated camera intrinsics and extrinsics, these points are back-projected into each image, where bilinear interpolation retrieves the corresponding feature vectors from all pyramid levels. Aggregated features are merged into the queries and fine-tuned using multi-head self-attention. This projection–refinement cycle repeats across multiple transformer decoder layers, with each stage producing updated classifications and 3D box regressions. DETR3D maintains an end-to-end pipeline that improves efficiency without sacrificing competitive detection accuracy by avoiding dense depth estimation, voxelization, and non-maximum suppression.
@@
[Xue_2025_CVPR] CorrBEV-fm 2025 CVPR Res101 Muiti-View 41.30 50.70 0.646 0.265 0.482 0.473 0.134
@@
[Xue_2025_CVPR] CorrBEV-sp 2025 CVPR Res101 Muiti-View 56.20 64.00 0.488 0.241 0.303 0.243 0.123
@@
[ji2025ropetr] RoPETR 2025 ArXiv V2-99 Muiti-View 52.90 61.40 0.537 0.255 0.289 0.229 0.195
@@
[li2024bevformer] BEVFormer 2024 IEEE Res101 Muiti-View 44.50 53.50 0.631 0.257 0.405 0.435 0.143
@@
[liu2024ray] RayDN 2024 ECCV Res50 Muiti-View 46.90 56.30 0.579 0.264 0.433 0.256 0.187
@@
[Wang_2023_ICCV] StreamPETR 2023 ICCV Res50 Muiti-View 45.00 55.00 0.569 0.267 0.413 0.265 0.196
@@
[Liu_2023_ICCV] PETRv2 2023 ICCV Res50 Muiti-View 34.90 45.60 0.700 0.275 0.580 0.437 0.187
@@
[li2023bevdepth] BEVDepth 2023 AAAI Res50 Muiti-View 35.10 47.50 0.639 0.267 0.479 0.428 0.198
@@
[Liu_2023_ICCV2] SparseBEV 2023 ICCV Res50 Muiti-View 44.80 55.80 0.581 0.271 0.373 0.247 0.190
@@
[Zong_2023_ICCV] HoP 2023 ICCV Res50 Muiti-View 40.00 51.00 0.607 0.275 0.515 0.286 0.216
@@
[jiang2023polarformer] PolarFormer 2023 AAAI Res101 Muiti-View 41.50 47.00 0.657 0.263 0.405 0.911 0.139
@@
[Wang_2023_CVPR] FrustumFormer 2023 CVPR Res101 Muiti-View 47.80 56.10 0.575 0.257 0.402 0.411 0.132
@@
[Li_2023_ICCV] FB-BEV 2023 ICCV V2-99 Muiti-View 53.70 62.40 0.439 0.250 0.358 0.270 0.128
@@
[Wang_2023_ICCV2] MV2D 2023 ICCV V2-99 Muiti-View 46.30 51.40 0.542 0.247 0.403 0.854 0.127
@@
[Xiong_2023_CVPR] CAPE 2023 CVPR V2-99 Muiti-View 52.50 61.00 0.503 0.242 0.361 0.306 0.114
@@
[Shu_2023_ICCV] 3DPPE 2023 ICCV VoV-99 Muiti-View 51.40 56.60 0.569 0.255 0.394 0.796 0.138
@@
[liu2022petr] PETR 2022 ECCV Res101 Muiti-View 37.00 45.50 0.647 0.251 0.433 0.933 0.143
@@
[li2022unifying] UVTR 2022 NeurIPS Res101 Muiti-View 37.90 48.30 0.731 0.267 0.350 0.510 0.200
@@
[li2023bevstereo] BEVStereo 2022 AAAI Res50 Muiti-View 37.20 50.00 0.597 0.270 0.438 0.367 0.190
@@
[wang2022probabilistic] PGD 2022 CoRL Res101 Muiti-View 38.60 44.80 0.626 0.245 0.451 1.509 0.127
@@
[wang2022detr3d] DETR3D 2022 CoRL Res101+FPN Muiti-View 41.20 47.90 0.641 0.255 0.394 0.845 0.133
@@
[chen2022graph] Graph-DETR3D 2022 ACM Res101 Muiti-View 41.80 47.20 0.668 0.250 0.440 0.876 0.139
@@
[Wang_2021_ICCV2] FCOS3D 2021 ICCV Res101 Muiti-View 35.80 42.80 0.690 0.249 0.452 1.434 0.124
@@
CorrBEV [Xue_2025_CVPR] improves upon transformer-based multi-view detection frameworks like BEVFormer. The task is completed by integrating a correlation-driven multi-modal approach to counteract the loss of discriminative information under occlusion. Its ”Multi-modal Prototype Generator” forms category-specific visual prototypes from cropped ground-truth object patches and derives complementary language prototypes from class names processed through a lightweight BERT encoder. These visual–linguistic embeddings are combined and introduced into the backbone features using a depth-wise correlation mechanism. The Correlation-guided Query Learner leverages these correlation features in two complementary ways: first, to initialize object queries using high-confidence spatial locations, improving recall for partially occluded objects; and second, to perform dual-path mixed sampling in which 3D queries aggregate both backbone and correlation-derived features for richer semantic–geometric representation. An occlusion-aware trainer further augments the system via pseudo-occlusion, randomly masking pixels in high-visibility samples, then a contrastive alignment loss that brings occluded and obvious instances of the same category closer in the embedding space. This modular design yields consistent gains across all occlusion levels and integrates seamlessly into diverse baseline architectures with minimal computational cost.
@@
Table 17 summarizes YOLO-based methods evaluated on the camera-based KITTI dataset for 2D object detection. Monocular 2D object detection methods have evolved to advance efficiency, representational depth, and domain-specific tuning. Early designs, such as YOLOV3-tiny [adarsh2020yolo], prioritized minimizing computation cost to improve efficiency with a precision trade-off. Later variants, including YOLOv5n [gao2025algorithm] with its scaled CSP-based backbone (C2f) derived from CSPDarknet53 and YOLOv6n [li2022yolov6] employing the EfficientRep backbone, coupled backbone revisions with refined neck structures and decoupled heads to raise accuracy without undermining real-time operation. YOLOv7 [wang2023yolov7] extended these gains by introducing re-parameterized convolution, a broadened trainable bag-of-freebies, and coarse-to-fine hierarchical label assignment. In turn, YOLOv8n [gao2025algorithm] replaced C3 modules with C2f blocks. It adopted an anchor-free head, simplifying the pipeline while preserving detection strength, whereas YOLOv8-Lite [yang2024yolov8] targeted resource-limited environments through compact backbone choices, lightweight feature aggregation, and attention mechanisms. ShuffYOLOX [he2023object] modified the YOLOX framework by utilizing the lightweight ShuffDet backbone and integrating ECA attention within the PAFPN, which reduces architectural complexity without negatively affecting detection capabilities.
@@
Domain-specific optimization has emerged as a critical trend within recent iterations. YOLO-Vehicle-v1s [li2024yolo] was designed for vehicle detection in adverse visual conditions, incorporating image dehazing and multimodal image-text fusion to enhance robustness. YOLOv8-RTDAV [gao2025algorithm] targeted road-target detection by embedding ECA-based C2f modules, DySample-based upsampling, and EIoU loss, yielding notable improvements in small-object accuracy. Transformer-CNN hybridization appeared in SwinYOLOv5s [an2025improved], which leveraged swin transformer attention with adaptive self-concat fusion to capture global context while mitigating false detections. Similarly, improved YOLOv5 [jia2023fast] employed NAS-pruned branches, coordinate attention, and structural re-parameterization to achieve both speed and high precision for small-object detection. The latest designs, YOLOv9t [wang2024yolov9] and YOLOv10n [wang2024yolov10], represent a shift toward more fundamental architectural changes: the introduction of the GELAN backbone and programmable gradient optimization in YOLOv9t improves training stability, while YOLOv10n adopts an NMS-free consistent dual-assignment strategy to harmonize efficiency with accuracy. Collectively, these methods illustrate a clear trajectory from minimalist, speed-oriented architectures to hybrid designs that balance inference efficiency with enhanced feature representation, addressing the increasingly diverse deployment requirements of modern detection systems.
@@
[an2025improved] SwinYOLOv5s 2025 MDPI MobileNetV3 Mono 96.00% 90.20% 95.70% 50.30M - -
@@
[gao2025algorithm] YOLOv8-RTDAV 2025 Scientific Reports C2f Mono 86.90% 81.10% 88.00% 2.81M 5.78MB 102.00